Temple Lunch Trucks, Inc. just paid a dividend of $3.50. Dividends are expected to grow at a rate of 4% per year from here on out. If the risk-free rate is 2.5%, the expected return on the market is 5% and Temple Lunch Trucks’ stock has twice the average market risk, what is the most that you should be willing to pay for a share of this stock today? please no excel or charts just straightforward calculations
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $3.50 * (1 + 4%) = $3.64

Required rate of return r can be calculated using the CAPM model, according to which

Required rate = Risk free rate + Beta * (Expected market return - Risk free rate)

Beta = 2 (since it is twice as risky as market)

Required rate = 2.5% + 2 * (5% - 2.5%)

Required rate = 2.5% + 5% = 7.5%

3.64 Po = 0.075 -0.04

Po = 104